Elimination: How do you solve these two problems:
r - 3p = 1
6r - p = 6

-2x + y = 3
x + 4y = 3

Answer by stanbon(75887)

r - 3p = 1
6r - p = 6
-------------
r = 3p+1
Substitute to solve for "p":
6(3p+1)-p = 6
17p= 0
p = 0
---------------------
Substitute in r=3p+1 to solve for "r":
r = 3*0+1 = 1
=======================================
-2x + y = 3
x + 4y = 3
--------
y = 2x+3
Substitute to solve for "x":
x + 4(2x+3) = 3
x + 8x+12 = 3
9x = -9
x = -1
------------
Cheers,
Stan H.